OpenStudy (swissgirl):

Let \( \mathbb{Z_8} \) = { 0,1,2,3,4,5,6,7,8} and \( \mathbb{Z_4} \) = {[0],[1],[2],[3]}. Define \( f: \mathbb{Z_8} --> \mathbb{Z_4}, g:\mathbb{Z_4} --> \mathbb{Z_8}, h:\mathbb{Z_8} --> \mathbb{Z_8}\) as follows: f(x)=[x+2], g([x])=2x and h(x)=2x+4

OpenStudy (anonymous):

Let \( \mathbb{Z_8} \) = { 0,1,2,3,4,5,6,7,8} and \( \mathbb{Z_4} \) = {[0],[1],[2],[3]}. Define \( f: \mathbb{Z_8} --> \mathbb{Z_4}, g:\mathbb{Z_4} --> \mathbb{Z_8}, h:\mathbb{Z_8} --> \mathbb{Z_8}\) as follows: f(x)=[x+2], g([x])=2x and h(x)=2x+4 and f(x)=[x+2]

OpenStudy (anonymous):

just so i could read it

OpenStudy (swissgirl):

Like all the numbers in \(\mathbb{Z_8} \) shld be having a line ontop if u know what i mean

OpenStudy (anonymous):

looks like \(f\) is there twice

OpenStudy (anonymous):

I guess it is looking for fog=h or something.

OpenStudy (anonymous):

yeah that is ok, but it is not clear what the question is

OpenStudy (anonymous):

just doing fog !

OpenStudy (swissgirl):

hahah yup i got so lost typing out all the extra info

OpenStudy (anonymous):

also i am confused by \(\mathbb{Z_8}\)

OpenStudy (swissgirl):

g(f(x))=h(x) so wld i have to show that its true for the 8 elements or whtvr

OpenStudy (anonymous):

i think it should be \(\{0,1,2,3,4,5,6,7\}\)

OpenStudy (anonymous):

and g(x)=[2x] , h(x)=[2x+4]

OpenStudy (anonymous):

ans satelite right that is Z8

OpenStudy (anonymous):

you can chase the diagram to do this, there are only a few elements to consider

OpenStudy (anonymous):

i.e. you can write \(f,g,h\) as sets of ordered pairs

OpenStudy (swissgirl):

omfggggggggg u r right

OpenStudy (anonymous):

i can write it if you like, but i am sure you can do it

OpenStudy (anonymous):